for the following equation, state the value of the discriminant and then describe the nature of the solutions: 10x^2 - 4x - 7 = 0
Answer by algebrahouse.com(1659)

The discriminant is b² - 4ac

If b² - 4ac = 0, there is one real number solution
If b² - 4ac < 0, there are two imaginary number solutions
If b² - 4ac > 0, there are two real number solutions

10x² - 4x - 7 = 0
a = 10, b = -4, and c = -7

b² - 4ac
= (-4)² - 4(10)(-7) {substituted values into the discriminant}
= 16 + 280 {multiplied}
= 296 {added}

296 > 0, there are two real number solutions
www.algebrahouse.com